# Do PAC-Learners Learn the Marginal Distribution?

### v2

@@ 2.5 Further Related Work
Our work and relaxed notions of density estimation also fit into a long line of literature studying learnability beyond uniform convergence. Most works in this area either study improved convergence rates via local complexities (e.g. local uniform convergence [42], local rademacher complexity [43]), or study models such as multi-class [20], private [44], or distribution learning [13] where uniform convergence and empirical risk minimization classically fail. Perhaps the most relevant to our setting is the study of Vapnik’s General Learning model [24], which also has no combinatorial characterization [45, 12], but can be characterized by the existence of stable learners [19]. We remark this characterization differs fairly substantially from our setting, not just in being distribution-free, but in the fact that stability is a property of a learning rule while we study connections between properties of the class and distribution family themselves.
3 PAC Learning vs. Intermediate Density Estimation
@@
so distances in the cover under $D$ change by at most $\varepsilon/2$ from their value under $D^{\prime}$. On the other hand, we are promised by the IDE that $h\in G_{T}$ with probability at least $41-\delta/$ 4. If both events occur, then $h$ is covered by $C_{T}$, meaning there exists $h^{\prime}\in C_{T}$ such that $d_{D}(h,h^{\prime})\leq 3\varepsilon/4$ as desired. ∎
We note that this observation generalizes the result of [21] that UBME is sufficient for PAC learning under finitely coverable distribution families as these have simple density estimators. With this out of the way we move to the more involved direction, showing intermediate density estimation fails necessity.
@@
This suggests the following strategy: after $O(\log\frac{1}{\delta})$ samples, simply output the majority label as a constant function. [^4] We note this can be made ‘proper’ simply by adding the all 00’s and all $1$’s function to the class. This only makes density estimation harder, so it does not effect the following lower bound. For any $\rho$ bounded away from $\frac{1}{2}$, a Chernoff bound promises that the learner picks the true majority label with probability at least $1-\delta$. Since the majority label has mass $1-\rho$, this gives the desired $(\rho,\delta)$-PAC-learner. ∎
@@
The crux of Proposition 3.4 lies in the fact that, given a few samples from a random $D_{x}^{\rho}$, while it may be easy to output some $x^{\prime}$ close to $x$ with high probability it is impossible to do so confidently. In other words, the learner should not be able to identify a large set of coordinates on which it knows it is correct due to the inherent noise in $D_{x}^{\rho}$. [^5] Algorithms with this type of guarantee are well-studied in the learning literature under a variety of names such as “Reliable and Probably Useful” [46], “Confident” [35], or “Knows What It Knows” [47]. We formalize this connection to confident learning via the following reduction.
@@
Noisy majority fails on any such input by construction, since it outputs majority label of a flipped coordinate. The probability there exists at least one coordinate that is flipped in each of $t$ samples is 1($161-(1-\rho^{t})^{n}\geq 15/$ 16 by our choice of constants. On the other hand, the expected number of coordinates with empirical density 00 or $1$ is at most $2n(1-\rho)^{t}\leq 2ne^{-\rho t}\leq\frac{n}{32}$ by our assumptions on $\rho$. By Markov’s inequality and a union bound both conditions hold with probability greater than $7/8$, so the algorithm cannot succeed with more than $1/8$ probability as desired. ∎
@@
On the other hand, it is easy to see there cannot be any bounded $(1/6,1/8)$-IDE for this class. For any fixed sample size $t$, the adversary can always choose $n$ such that $t_{n}>t$ (since $\{t_{n}\}$ is increasing), and corresponding $\rho_{n}$ satisfying the conditions of Proposition 3.4. Since $\mathcal{H}_{n}^{\rho_{n}}$ then has no $(1/6,1/8)$-IED on $t$ samples, this completes the proof. ∎
@@
4 PAC Learning vs. Weak Density Estimation
In this section, we show that weak density estimation is equivalent to PAC Learning in the distribution-free setting, and necessary but not sufficient in the distribution-family setting. We start with the distribution-free case. We note that necessity follows from the classical necessity of density estimation, so it is sufficient to prove sufficiency. To this end, we show the existence of a WDE implies finite VC dimension.
Proposition 4.1.
Any $(\frac{1}{4},\frac{1}{4})$-WDE for a class $(X,H)$ requires at least
$n_{WDE}\left(1/4,1/4\right)\geq\Omega(\sqrt{VC(H)})$
samples.
Proof.
Denote $d\coloneqq VC(H)$. We will rely on the fact that it is impossible for any algorithm on $t=o(\sqrt{d})$ samples to distinguish between the following distributions over $\mathbb{F}_{2}^{d}$:
With this in mind, it is sufficient to show that a WDE for $(X,H)$ on $o(\sqrt{d})$ samples gives such a distinguisher. ∎
We now move to necessity in the distribution-family model.
@@
Proof.
The proof appeals to the non-uniform covering technique of HKLM [18]. In particular, draw an unlabeled sample $T$ of size $n_{\text{PAC}}(\varepsilon/3,\delta)$ and run a PAC-learner $\mathcal{L}$ across all possible labelings of $T$ to generate the set
$C_{T}\coloneqq\{\mathcal{L}(T,h(T)):\leavevmode\nobreak\ h\in H\}.$
HKLM [18, Lemma 5.2] observed $C_{T}$ non-uniformly covers $H$ in the following sense:
@@
Formally, draw a fresh sample $T^{\prime}$ of $O(\frac{\log|C_{T}|+\log(1/\delta)}{\varepsilon^{2}})$ unlabeled examples. By Chernoff and Union bounds, the empirical estimator $\hat{d}_{T^{\prime}}$ is close to the true distance metric $d_{D}$ with high probability across elements of $C_{T}$:
@@
$G_{(T,T^{\prime}),D}\coloneqq\{g\in H:\leavevmode\nobreak\ d_{D}(g,c_{T}(g))\leq\varepsilon/3\}.$
@@
Putting everything together, we have that each $h\in H$ lies in $G_{(T,T^{\prime}),D}$ with probability at least $1-\delta$, and that all estimates in $G_{(T,T^{\prime}),D}$ are $\varepsilon$-accurate with probability at least $1-\delta$ giving the desired WDE. ∎

### v3

@@ 2.3 Further Related Work
Learning Beyond Uniform Convergence. Our work also fits into a long line of literature studying learnability beyond uniform convergence, either through notions of local complexity (e.g. Panchenko (2002); Bartlett et al. (2005)) or models where uniform convergence and empirical risk minimization fail completely (e.g. Shalev-Shwartz et al. (2010); David et al. (2016)). The former works relate qualitatively to our weakened notions of estimation which only require good estimates for a subset of hypotheses in the class. Most relevant, however, is work in the latter line of Shalev-Shwartz et al. (2010), who characterize Vapnik’s General Learning model (which fails uniform convergence and has no combinatorial characterization (Ben-David et al., 2019)) via stability. Their work differs from our setting both in that the model is distribution-free, and that stability is a property of a learning rule while we study direct connections between the class and distribution family themselves.
3 Results and Proof Overview
We now overview our main results and sketch their proofs. In Section 3.1 we show IDE is sufficient but not necessary for learning. In Section 3.2 we show WDE is necessary but not sufficient. In Section 3.3 we overview the connections between WDE, UBME, and Uniform Estimation.
3.1 PAC Learning vs. Intermediate Density Estimation
In this section we study the relation between PAC Learning and intermediate density estimation. All proofs are sketched or otherwise shortened, and detailed versions can be found in Section 4. We start with the simple observation that (under UBME), IDE is sufficient to PAC-learn.
Proposition 3.1.
Let $(\mathscr{P},X,H)$ be any class with metric entropy $m(\varepsilon)$ and an intermediate density estimator on $n_{IDE}(\varepsilon,\delta)$ samples. Then $(\mathscr{P},X,H)$ has a PAC learner satisfying
$n_{\text{PAC}}(\varepsilon,\delta)\leq O\left(n_{IDE}(\varepsilon/4,\delta/4)+\frac{\log(m(\varepsilon/4))+\log(1/\delta)}{\varepsilon}\right).$
Proof 3.2.
Run the IDE on an unlabeled sample $T$ of size $n_{\text{IDE}}(\varepsilon/4,\delta/4)$ to get the approximate metric $\tilde{d}_{T}$ and set of well-estimated functions $G_{T}$. We are promised that with probability $41-\delta/$ 4:
Assume this is the case. Because $G_{T}$ is known, we can find $D^{\prime}\in\mathscr{P}$ which realizes these estimates on $G_{T}$ up to an error of $\varepsilon/4$, and therefore satisfies $TV_{G_{T}\Delta G_{T}}(D,D^{\prime})\leq\varepsilon/2$. Since $(\mathscr{P},H)$ has UBME, there exists an $\frac{\varepsilon}{4}$-cover $C_{T}$ of $G_{T}$ under $D^{\prime}$ of size at most $m(\varepsilon/4)$. It is enough to argue that with probability at least $21-\delta/$ 2, there exists a hypothesis in $C_{T}$ with true risk at most $\frac{3\varepsilon}{4}$. In this case, if we simply draw $O(\frac{\log|C_{T}|+\log(1/\delta)}{\varepsilon})$ fresh labeled samples and output the hypothesis with lowest empirical error, a Chernoff and Union bound imply the resulting hypothesis has error at most $\varepsilon$ with probability at least $41-\delta/$ 4. By a union bound, all steps in this process succeed simultaneously except with probability $\delta$, which gives the desired learner.
It is left to show a good hypothesis exists in $C_{T}$ with high probability. To see this, observe that if $TV_{G_{T}\Delta G_{T}}(D,D^{\prime})\leq\varepsilon/2$ (which occurs except with probability $\delta/4$), $C_{T}$ is a $43\varepsilon/$ 4-cover for $G_{T}$ under the true distribution $D$. This follows simply from noting that for any $g,g^{\prime}\in G_{T}$ we have:
$d_{D}(g,g^{\prime})$ $=d_{D}(g,g^{\prime})-d_{D^{\prime}}(g,g^{\prime})+d_{D^{\prime}}(g,g^{\prime})\leq TV_{G_{T}\Delta G_{T}}(D,D^{\prime})+d_{D^{\prime}}(g,g^{\prime}),$
so distances in the cover under $D$ change by at most $\varepsilon/2$ from their value under $D^{\prime}$. On the other hand, we are promised by the IDE that $h\in G_{T}$ with probability at least $41-\delta/$ 4. If both events occur, $h$ is covered by $C_{T}$, meaning $\exists h^{\prime}\in C_{T}$ such that $d_{D}(h,h^{\prime})\leq 3\varepsilon/4$ as desired.
We note this generalizes the result of Kulkarni and Vidyasagar (1997) that UBME is sufficient for PAC learning under finitely coverable distribution families as these have simple density estimators. We now show the more involved direction: intermediate density estimation fails necessity.
Theorem 3.3.
There exists a class $(\mathscr{P},X,H)$ which is PAC-learnable in $n_{\text{PAC}}(\varepsilon,\delta)\leq\tilde{O}\left(\frac{\log(1/\delta)}{\varepsilon^{2}}\right)$ samples, but has no intermediate density estimator.
We overview the construction, which is based on dictators on the noisy cube. Let $X_{n}=\{0,1\}^{n}$ denote the $n$-dimensional hypercube, and for $x\in X_{n}$ write $x_{i}$ to denote the $i$th coordinate of $x$. Our hypothesis class $H_{n}\coloneqq\{\mathbf{1}_{i}:i\in[n]\}$ will consist of the set of dictators on the cube:
For every $x\in\{0,1\}^{n}$ and $\rho>0$, let $D_{x}^{\rho}$ denote the distribution that samples $y\in\{0,1\}^{n}$ by independently setting each bit $y_{i}$ to $x_{i}$ with probability $1-\rho$, and otherwise to $1-x_{i}$. Let $\mathscr{P}_{n}^{\rho}\coloneqq\{D^{\rho}_{x}\}_{x\in X_{n}}$ be the family of all such distributions. We argue $(\mathscr{P}_{n}^{\rho},X_{n},H_{n})$ is learnable up to error $\rho$ in $O_{\delta}(1)$ samples, while any IDE requires $\Omega_{\rho}(\log(n))$ samples. The result then follows from taking an infinite disjoint union with $n\to\infty$ and $\rho\to 0$. We focus here on fixed $n,\rho$ and give the full construction in Section 4. We’ll start with learnability, which simply stems from the fact that our class is ‘almost constant,’ and can therefore be easily approximated by a constant function.
Lemma 3.4.
Proof 3.5.
This suggests the following strategy: after $O(\log\frac{1}{\delta})$ samples, output the majority label as a constant function. [^4] We note this can be made ‘proper’ simply by adding the all 00’s and all $1$’s function to the class. This only makes IDE harder, so it does not effect the following lower bound. For $\rho\leq 1/6$, Chernoff promises the learner picks the true majority with probability at least $1-\delta$. Since the majority label has mass $1-\rho$, this gives the desired $(\rho,\delta)$-PAC-learner.
Proposition 3.6.
For all $n>2$ and $\frac{1}{6}>\rho>\frac{12\log(1/\rho)}{\log(n)}$, any IDE for $(\mathscr{P}_{n}^{\rho},X_{n},H_{n})$ satisfies
$n_{\text{IDE}}(1/6,1/8)\geq\Omega\left(\frac{\log(n)}{\log(1/\rho)}\right).$
Formally, it is convenient to reduce Proposition 3.6 to showing no algorithm can ‘confidently’ compute even a portion of the string $x$ underlying some $D_{x}^{\rho}\in\mathscr{P}_{n}^{\rho}$ in the following sense.
Lemma 3.7.
  1. 1.
    $\hat{x}$ is $\bot$’ on at most $n/2$ coordinates,
  2. 2.
    $\hat{x}$ agrees with $x$ otherwise.
The proof, which follows from observing that $\tilde{d}_{T}$ can be used to compute $x_{i}$ whenever $\mathbf{1}_{i}\in G_{T}$, is deferred to Section 4 (Lemma 4.7). It is now enough to show no such learner exists.
Proof 3.8 (Sketch, Proposition 3.6).
By Lemma 3.7 and Yao’s minimax principle, it is enough to exhibit a distribution over $\mathscr{P}_{n}^{\rho}$ such that any deterministic confident learner on $t$ samples fails with probability at least $7/8$. We will use the uniform distribution over $\mathscr{P}^{\rho}_{n}$, which corresponds to choosing $D_{x}^{\rho}$ for a uniformly random $x\in\{0,1\}^{n}$. Observe it is sufficient to consider learners $\mathcal{A}$ that output $\bot$ on the noisiest $\frac{n}{2}$ coordinates [^5] I.e. coordinates with empirical density closest to $1/2$, ties broken arbitrarily. and output the majority bit otherwise, since these maximize the learner’s success rate (see Proposition 4.6 for details). The idea is to then observe that unless $t$ is sufficiently large, for any choice of $D_{x}^{\rho}$ the following two events almost certainly occur across $T\sim(D_{x}^{\rho})^{t}$:
  1. 1.
    There is a coordinate $i\in[n]$ flipped by noise in every example: $\forall y\in T:y_{i}\neq x_{i}$
  2. 2.
    There are at least $n/2$ coordinates $i\in[n]$ with empirical density $10<\frac{1}{t}\sum_{y\in T}y_{i}<$ 1.
$\mathcal{A}$ fails on any such input by construction, since it outputs the empirical majority on a flipped coordinate. The probability that there exists at least one coordinate that is flipped in each of $t$ samples is 1($161-(1-\rho^{t})^{n}\geq 15/$ 16 by our choice of constants. On the other hand, the expected number of coordinates with empirical density 00 or $1$ is at most $2n(1-\rho)^{t}\leq 2ne^{-\rho t}\leq\frac{n}{32}$ by our assumptions on $\rho$. By Markov’s inequality and a union bound both conditions hold with probability greater than $7/8$, so the algorithm cannot succeed with more than $1/8$ probability as desired.
3.2 PAC Learning vs. Weak Density Estimation
We now show weak density estimation is necessary but not sufficient to PAC-learn. All proofs are sketched or otherwise shortened, and detailed versions can be found in Section 5.
Proposition 3.9.
Let $(\mathscr{P},X,H)$ be a class that is PAC-learnable in $n_{\text{PAC}}(\varepsilon,\delta)$ samples. Then $(\mathscr{P},X,H)$ has a weak density estimator satisfying
$n_{\text{WDE}}(\varepsilon,\delta)\leq O\left(n_{\text{PAC}}(\varepsilon/3,\delta)+\frac{\log(\Pi_{H}(n_{\text{PAC}}(\varepsilon/3,\delta)))+\log(1/\delta)}{\varepsilon^{2}}\right),$
where we recall $\Pi_{H}$ is the growth function $\Pi_{H}(n)\coloneqq\max_{S\subset X:|S|=n}|\{H|_{S}\}|$.
Proof 3.10.
Run a PAC-learner $\mathcal{L}$ across all labelings of a sample $T$ of size $n_{\text{PAC}}(\varepsilon/3,\delta)$ to generate
$C_{T}\coloneqq\{\mathcal{L}(T,h(T)):~{}h\in H\}.$
Hopkins et al. (2024) observed $C_{T}$ is a ‘non-uniform cover’ in the sense that for each $h\in H$, if we define $c_{T}:H\to C_{T}$ by $c_{T}(h)\coloneqq\mathcal{L}(T,h(T)),$ then $c_{T}(h)$ is close to $h$ with high probability:
$\forall h\in H:\Pr_{T}[d_{D}(h,c_{T}(h))>\varepsilon/3]<\delta.$
To build our WDE, we directly estimate distances on $C_{T}$ and extend to $H$ via $c_{T}$. Draw a fresh sample $T^{\prime}$ of $O(\frac{\log|C_{T}|+\log(1/\delta)}{\varepsilon^{2}})$ unlabeled examples. By Chernoff and Union bounds, the empirical estimator $\hat{d}_{T^{\prime}}(h,h^{\prime})\coloneqq\frac{1}{|T^{\prime}|}\sum\limits_{x\in T^{\prime}}\mathbf{1}\{h(x)\neq h^{\prime}(x)\}$ is close to $d_{D}$ with high probability on $C_{T}$:
$\tilde{d}_{T\cup T^{\prime}}(h,h^{\prime})\coloneqq\hat{d}_{T^{\prime}}(c_{T}(h),c_{T}(h^{\prime})).$
We argue this estimate is accurate across all hypotheses which are close to their image under $c_{T}$:
$G_{T\cup T^{\prime},D}\coloneqq\{g\in H:~{}d_{D}(g,c_{T}(g))\leq\varepsilon/3\}.$
Assuming our empirical estimates on $C_{T}$ are good, note for any $g,g^{\prime}\in G_{T\cup T^{\prime},D}$ we can write:
$|\tilde{d}_{T\cup T^{\prime}}(g,g^{\prime})-d_{D}(g,g^{\prime})|$ $=|\hat{d}_{T^{\prime}}(c_{T}(g),c_{T}(g^{\prime}))-d_{D}(g,g^{\prime})|$
$\leq\varepsilon,$
Thus we have that each $h\in H$ lies in $G_{T\cup T^{\prime},D}$ with probability at least $1-\delta$, and all estimates in $G_{T\cup T^{\prime},D}$ are $\varepsilon$-accurate with probability at least $1-\delta$, which gives the desired WDE.
Given the closeness of weak and intermediate density estimation, one might hope to amplify weak estimation to a PAC-learner. Unfortunately, the main barrier to such a proof (lacking knowledge of the well-estimated set) is inherent, and we are able to give a very strong refutation of WDE’s sufficiency: even a perfect weak estimator (i.e. one with error $\varepsilon=0$) isn’t sufficient to PAC-learn.
Theorem 3.11.
There exists a class $(\mathscr{P},X,H)$ with metric entropy $m(\varepsilon)\leq O(2^{1/\varepsilon})$ and a WDE on $n_{\text{WDE}}(0,\delta)\leq O\left(\log(1/\delta)\right)$ samples which is not PAC-learnable.
We first describe the construction, which is similar in spirit to the UBME counterexample of Dudley et al. (1994). Let $X=\{0,1,2\}^{\mathbb{N}}$. For each $i\in\mathbb{N}$, let $\varepsilon_{i}\coloneqq\frac{1}{\log(i+255)}$, define $f_{i}\coloneqq\mathbf{1}[x_{i}=0]$, and $D_{i}$ to be the distribution over $x\in X$ where $x_{i}$ is chosen from $\{0,1\}$ uniformly at random and all remaining coordinates $j\neq i$ are drawn independently from the categorical distribution
$\Pr[x_{j}=0]=\Pr[x_{j}=1]=\epsilon_{j},~{}\text{and }\Pr[x_{j}=2]=1-2\epsilon_{j}.$
Proposition 3.12.
$(\mathscr{P},X,H)$ has a weak density estimator on $n_{\text{WDE}}(0,\delta)\leq O(\log(1/\delta))$ samples.
Proof 3.13.
Fix any $f_{j},f_{k}$, and observe that across the choices of $D_{i}$, $d_{D_{i}}(f_{j},f_{k})$ depends only on whether $i\in\{j,k\}$. Namely, for any $j\neq i$ we have $d_{D_{i}}(f_{i},f_{j})=\frac{1}{2}$, while any $i\neq j\neq k$ satisfy $d_{D_{i}}(f_{j},f_{k})=\epsilon_{j}(1-\epsilon_{k})+\epsilon_{k}(1-\epsilon_{j})$. Thus we can compute distances simply by checking if $i\in\{j,k\}$. With this in mind, given a sample $T$ of size $O(\log(1/\delta))$, let $\mu_{\ell}^{T}\coloneqq\frac{1}{t}\sum\limits_{y\in T}y_{\ell}$ and define:
$\tilde{d}_{T}(f_{j},f_{k})=\begin{cases}\epsilon_{j}(1-\epsilon_{k})+\epsilon_{k}(1-\epsilon_{j})&\text{if }\mu_{j}^{T},\mu_{k}^{T}<1/3\\ 1/2&\text{otherwise,}\end{cases}$
and let $G_{T,D}$ be the set of functions with good empirical averages:
[…]
Fix any $D_{i}\in\mathscr{P}$ and $f_{j}\in H$. First, observe that by a Chernoff bound, $f_{j}\in G_{T,D_{i}}$ with probability at least $1-\delta$. Second, notice that for all functions $f_{\ell}\in G_{T,D_{i}}$, we are assured that $\mu_{\ell}^{T}<1/3$ if and only if $\ell\neq i$ (since $\mathbb{E}[f_{i}]=1/2$, while $\mathbb{E}[f_{\ell}]\leq 1/8$ for $\ell\neq i$). Since this condition exactly determines the distances between hypotheses, $\tilde{d}_{T}$ has zero error across $G_{T,D_{i}}$ as desired.
Next, we claim $(\mathscr{P},X,H)$ has small metric entropy, which simply follows from observing that for any fixed $D_{i}$, most hypotheses are close to the all 00’s function (see Section 5, Proposition 5.6).
Proposition 3.14.
$(\mathscr{P},X,H)$ has metric entropy at most $m(\varepsilon)\leq O(2^{1/\varepsilon})$.
Finally, we argue our class is not PAC-learnable. The proof has a similar flavor to our lower bound against IDEs: the learner cannot distinguish between ground truth and hypotheses that happen to look consistent due to a small amount of noise (here generated by the $\epsilon_{i}$’s).
Proposition 3.15.
$(\mathscr{P},X,H)$ is not PAC-learnable.
Proof 3.16 (Sketch).
By Yao’s principle it suffices to find for every $t\in\mathbb{N}$ a randomized strategy for the adversary over which any deterministic learner on $t$ samples fails with at least constant probability. For each $n\in\mathbb{N}$, consider the strategy that selects from the set $\{D_{i},f_{i}\}_{i\in[n]}$ with probability $\Pr[(D_{i},f_{i})]\propto\epsilon_{i}^{t}$. For each $j\in[n]$, call a sample $S$ consistent with $(D_{j},f_{j})$ if for all $(x,y)\in S$, we have $x\in\text{supp}(D_{j})$ and $f_{j}(x)=y$. The idea is to argue that when the learner’s sample is consistent with multiple coordinates, the posterior choices over consistent $(D_{j},f_{j})$ are uniform and the learner cannot distinguish the ground truth. On the other hand, for large enough $n\gg t$, there are almost always multiple consistent coordinates. See Section 5 (Proposition 5.8) for further details.
3.3 Density Estimation, Uniform Convergence, and UBME
Finally, we compare density estimation with prior paradigms in the literature. All proofs are sketched or otherwise shortened, and detailed versions can be found in Section 6. We start by introducing a relaxed variant of uniform convergence called uniform estimation.
Definition 3.17 (Uniform Estimation).
We say $(\mathscr{P},X,H)$ satisfies the uniform estimation property if there exist a family of estimators $\{\mathcal{E}_{S}\}_{S\in(X\times\{0,1\})^{*}}:H\to\mathbb{R}_{+}$ and a function $n=n_{\text{UE}}(\varepsilon,\delta)$ such that for all $\varepsilon,\delta>0$, $D\in\mathscr{P}$, and $h\in H$:
Uniform estimation is the natural supervised analog of density estimation, estimating $err_{D\times h}(h^{\prime})=d_{D}(h,h^{\prime})$ from labeled samples. We formalize this connection by showing the two are equivalent so long as the underlying class has UBME. [^6] Note unlike the distribution-free case this is trivially false otherwise since uniform estimation implies UBME, but any triple $(D,X,H)$ has a trivial density estimator.
Theorem 3.18.
If $(\mathscr{P},X,H)$ has a uniform estimator on $n_{\text{UE}}(\varepsilon,\delta)$ samples, then $(\mathscr{P},X,H)$ has a density estimator satisfying
where $\varepsilon^{\prime}=O(\varepsilon)$, $\delta^{\prime}=O(\frac{\delta}{\Pi_{H}(n_{UE}(\varepsilon/4,\delta/2))})$. Conversely, if $(\mathscr{P},X,H)$ has a density estimator on $n_{\text{DE}}(\varepsilon,\delta)$ samples and has metric entropy $m(\varepsilon)$, then $(\mathscr{P},X,H)$ has a uniform estimator satisfying
$n_{\text{UE}}(\varepsilon,\delta)\leq O\left(n_{\text{DE}}(\varepsilon/4,\delta/2)+\frac{\log(m(\varepsilon/4))+\log(1/\delta)}{\varepsilon^{2}}\right).$
Theorem 3.18 implicitly revolves around the fact that under UBME, both uniform and density estimation are equivalent to building a (bounded) uniform $\varepsilon$-cover $C$ and a corresponding covering map $c:H\to C$ such that $\forall h\in H:d_{D}(c(h),h)\leq\varepsilon.$ This type of structure allows us to apply Yatracos (1985)’s trick extending empirical estimates on $C$ to good estimates on all of $H$. We now overview the constructions underlying Theorem 3.18, leaving proof details to Section 6 (Theorem 6.1).
DE + UBME $\implies$ UE:
Running the UE outputs a distribution $D^{\prime}$ s.t. $TV_{H\Delta H}(D^{\prime},D)$ is small. UBME promises that $D^{\prime}$ has a known cover $C$ and covering map $c$, which is in turn a covering map for $D$ as well (as in Proposition 3.1). We can then define our estimators by drawing a labeled sample $S$, and extending the empirical error of $S$ on $C$ to all of $H$ via $\mathcal{E}_{S}(h)\coloneqq\text{err}_{S}(c(h))$.
UE $\implies$ DE:
Hopkins et al. (2024) showed it is possible to build a cover $C$ for any learnable class from unlabeled samples (see Section 6, Lemma 6.3). We show that using the uniform estimator, it is also possible to construct $C$’s associated covering map. [^7] Note this is not in general possible for any learnable class, as this would refute Theorem 3.3. Draw an unlabeled sample $T$, and run the uniform estimator across every possible labeling of $T$ by $C$ to obtain $\{\mathcal{E}_{(T,h(T))}\}_{h\in C}$. We define a covering map by sending $h\in H$ to its closest element in $C$ according to these estimates: $c(h)\coloneqq\text{argmin}_{h^{\prime}\in C}\ \mathcal{E}_{(T,h^{\prime}(T))}(h)$. Now draw an unlabeled sample $T$, define $\tilde{d}_{T}$ by extending empirical estimates on $C$ to $H$ as $\tilde{d}_{T^{\prime}}(h,h^{\prime})\coloneqq\hat{d}_{T^{\prime}}(c(h),c(h^{\prime}))$, and output any distribution $D^{\prime}\in\mathscr{P}$ approximately satisfying all estimates.
It is left to examine the connection between density estimation and UBME. We show the latter actually directly implies weak density estimation.
Proposition 3.19.
Any class $(\mathscr{P},X,H)$ with metric entropy $m(\varepsilon)$ has a WDE satisfying
$n_{\text{WTV}}(\varepsilon,\delta)\leq O\left(\frac{\log(m(\varepsilon/8))+\log(1/\delta)}{\varepsilon^{2}}\right).$
Proof 3.20 (Sketch).
We argue it is sufficient to simply output the empirical distance estimator on a large enough unlabeled sample $T$. UBME promises the existence of an (unknown) cover $C$ and covering map $c$. Let $G_{T,D}$ be the set of all $h\in H$ s.t. $d_{D}(h,c(h))$ is estimated to within $O(\varepsilon)$ error. The idea is to observe that the empirical distance of any two $h,h^{\prime}\in G_{T,D}$ can be factorized through their distances to $c(h)$ and $c(h^{\prime})$ without any knowledge of $c$, since:
$|\hat{d}_{T}(h,h^{\prime})-d_{D}(h,h^{\prime})|$ $\leq|\hat{d}_{T}(c(h),c(h^{\prime}))-d_{D}(c(h),c(h^{\prime}))|+O(\varepsilon).$
Thus as long as we draw a sufficient number of samples to estimate distances on $C$, just the existence of this cover is sufficient to imply $\hat{d}_{T}$ is a WDE. See Section 6 (Proposition 6.9) for details.
Note that while Proposition 3.19 subsumes Proposition 3.9 since any PAC-learnable class has UBME (Benedek and Itai, 1991), they give quantitatively different sample complexity guarantees. Combined with Dudley et al. (1994), Proposition 3.19 also implies insufficiency of WDEs, but the approach is inherently lossy and cannot handle the ‘perfect’ $\varepsilon=0$ case Theorem 3.11 covers.
Acknowledgements
SL was supported by a Simons Investigator Award (##\##929894). DK was supported by by NSF Medium Award CCF-2107547 and NSF Award CCF-1553288 (CAREER).
[…]
@@ Acknowledgements
so distances in the cover under $D$ change by at most $\varepsilon/2$ from their value under $D^{\prime}$. On the other hand, we are promised by the IDE that $h\in G_{T}$ with probability at least $41-\delta/$ 4. If both events occur, then $h$ is covered by $C_{T}$, meaning there exists $h^{\prime}\in C_{T}$ such that $d_{D}(h,h^{\prime})\leq 3\varepsilon/4$ as desired.
We note that this observation generalizes the result of Kulkarni and Vidyasagar (1997) that UBME is sufficient for PAC learning under finitely coverable distribution families as these have simple density estimators. With this out of the way we move to the more involved direction, showing intermediate density estimation fails necessity.
@@
This suggests the following strategy: after $O(\log\frac{1}{\delta})$ samples, simply output the majority label as a constant function. [^8] We note this can be made ‘proper’ simply by adding the all 00’s and all $1$’s function to the class. This only makes density estimation harder, so it does not effect the following lower bound. For any $\rho$ bounded away from $\frac{1}{2}$, a Chernoff bound promises that the learner picks the true majority label with probability at least $1-\delta$. Since the majority label has mass $1-\rho$, this gives the desired $(\rho,\delta)$-PAC-learner.
@@
The crux of Proposition 4.6 lies in the fact that, given a few samples from a random $D_{x}^{\rho}$, while it may be easy to output some $x^{\prime}$ close to $x$ with high probability it is impossible to do so confidently. In other words, the learner should not be able to identify a large set of coordinates on which it knows it is correct due to the inherent noise in $D_{x}^{\rho}$. [^9] Algorithms with this type of guarantee are well-studied in the learning literature under a variety of names such as “Reliable and Probably Useful” (Rivest and Sloan, 1988), “Confident” (Kane et al., 2017), or “Knows What It Knows” (Li et al., 2001). We formalize this connection to confident learning via the following reduction.
@@
Noisy majority fails on any such input by construction, since it outputs majority label of a flipped coordinate. The probability there exists at least one coordinate that is flipped in each of $t$ samples is 1($161-(1-\rho^{t})^{n}\geq 15/$ 16 by our choice of constants. On the other hand, the expected number of coordinates with empirical density 00 or $1$ is at most $2n(1-\rho)^{t}\leq 2ne^{-\rho t}\leq\frac{n}{32}$ by our assumptions on $\rho$. By Markov’s inequality and a union bound both conditions hold with probability greater than $7/8$, so the algorithm cannot succeed with more than $1/8$ probability as desired.
@@
On the other hand, it is easy to see there cannot be any bounded $(1/6,1/8)$-IDE for this class. For any fixed sample size $t$, the adversary can always choose $n$ such that $t_{n}>t$ (since $\{t_{n}\}$ is increasing), and corresponding $\rho_{n}$ satisfying the conditions of Proposition 4.6. Since $\mathcal{H}_{n}^{\rho_{n}}$ then has no $(1/6,1/8)$-IED on $t$ samples, this completes the proof.
@@
[…]
In this section, we show that weak density estimation is necessary but not sufficient to PAC-learn. We start with necessity.
@@
Proof 5.2.
The proof appeals to the non-uniform covering technique of Hopkins et al. (2024). In particular, draw an unlabeled sample $T$ of size $n_{\text{PAC}}(\varepsilon/3,\delta)$ and run a PAC-learner $\mathcal{L}$ across all possible labelings of $T$ to generate the set
$C_{T}\coloneqq\{\mathcal{L}(T,h(T)):~{}h\in H\}.$
HKLM (Hopkins et al., 2024, Lemma 5.2) observed $C_{T}$ non-uniformly covers $H$ in the following sense:
@@
Formally, draw a fresh sample $T^{\prime}$ of $O(\frac{\log|C_{T}|+\log(1/\delta)}{\varepsilon^{2}})$ unlabeled examples. By Chernoff and Union bounds, the empirical estimator $\hat{d}_{T}(h,h^{\prime})\coloneqq\frac{1}{|T|}\sum\limits_{x\in T}\mathbf{1}\{h(x)\neq h^{\prime}(x)\}$ is close to the true distance metric $d_{D}$ with high probability across elements of $C_{T}$:
@@
$G_{(T,T^{\prime}),D}\coloneqq\{g\in H:~{}d_{D}(g,c_{T}(g))\leq\varepsilon/3\}.$
@@
Putting everything together, we have that each $h\in H$ lies in $G_{(T,T^{\prime}),D}$ with probability at least $1-\delta$, and that all estimates in $G_{(T,T^{\prime}),D}$ are $\varepsilon$-accurate with probability at least $1-\delta$ giving the desired WDE.